Commute times in the U.S. are heavily skewed to the right. We select a random sample of 200 people from the 2000 U.S. Census who reported a non-zero commute time.

In this sample the mean commute time is 27.5 minutes with a standard deviation of 18.9 minutes. Can we conclude from this data that the mean commute time in the U.S. is less than half an hour? Conduct a hypothesis test at the 5% level of significance.

What is the p-value for this hypothesis test?

Your answer should be rounded to 4 decimal places.
Answer #1

Solution :

\mu= 30

\bar x =27.5

\sigma =18.9

n = 200

This is the less tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 30

Ha : \mu < 30

Test statistic = z

= (\bar x - \mu ) / \sigma / \sqrt n

= (27.5-30) / 18.9/ \sqrt 200

= -1.87

P(z <-1.87 ) = 0.0307

P-value = 0.0307

\alpha = 0.05  

P-value < \alpha

Reject the null hypothesis .